OpenStudy (anonymous):

HELP PLEASE What is the range of the following function? y=1/x+2 -1 A.{y:y ER, y ≠ -1} B. {y:y ER, y ≠ 2} C. {y:y ER} D. {y:y ER, y ≠ 1}

OpenStudy (anonymous):

what is ER?

OpenStudy (anonymous):

this

OpenStudy (anonymous):

ooooh i see \(y\in \mathbb{R}\) i get it !

OpenStudy (anonymous):

how do you do it?

OpenStudy (anonymous):

is it \[f(x)=\frac{1}{x+2}-1\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The range of a function is where the y-coordinate exists. Is the equation \[y =\frac{ 1 }{ x+2 } - 1\]?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

then we can think it is pretty clear that \[\frac{1}{x+2}\] is never zero, because the numerator is 1

OpenStudy (anonymous):

and since \[\frac{1}{x+2}\] is never zero, then \[\frac{1}{x+2}-1\] is never \( -1\)

OpenStudy (anonymous):

so A?

OpenStudy (anonymous):

so i would go with \[\{y:y\in \mathbb{R}, y\neq -1\}\]

OpenStudy (anonymous):

which i have to say is a rather weird way to write this we pretty much assume that \(y\) is a real number to start with

OpenStudy (anonymous):

Thanks it was :)

OpenStudy (anonymous):

to answer, yes, A

OpenStudy (anonymous):

For the record, satellite, that is the correct way to write it on all AP Math tests.

OpenStudy (anonymous):

why not \[\{y\in \mathbb{R}:y\neq -1\}\] that looks at least half way normal